### v1

@@ 3 Balanced Sparse-Cut via spectral Cut-Matching
In this section we state our main theorem and present our algorithm, whose analysis appears in the next section. The result of [2] for finding a balance sparse cut for the case of conductance is a special case of our theorem in which the vertex measure corresponds to degrees (i.e. $\mu(v)=deg_{G}(v)$ for every $v\in V$).
@@
Given a graph $G=(V,E)$, a vertex measure $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$ that satisfies Assumption 2.1 and a parameter $\phi>0$, there exists a randomized algorithm which takes $\tilde{O}\left(m\right)$ time [^4] The $\tilde{O}$ hides $\log^{c}n$ factors as well as $\log^{c}\mu^{\star}$ factor, where $\mu^{\star}=\frac{\max(\mu(\cdot))}{\min(\mu(\cdot))}$. and must end in one of the following three cases:
@@
    We certify that $G$ has $\mu$-expansion $\Phi^{\mu}(G)=\Omega(\phi)$ with high probability.
@@
    We find a cut $(R,A)$ in $G$ of $\mu$-expansion $\Phi^{\mu}_{G}(R,A)=O(\phi\log n)$, and $\mu(R),\mu(A)$ are both $\Omega(\frac{\mu(V)}{\log n})$, i.e, we find a relatively balanced low conductance cut.
@@
    We find a cut $(R,A)$ with $\Phi^{\mu}_{G}(R,A)\leq c_{0}\phi\log n$ for some constant $c_{0}$, and $\mu(R)\leq\frac{\mu(V)}{10c_{0}\log n}$, and with high probability $A$ is a near $(\Omega(\phi),\mu)$-expander in $G$.
This theorem is a straightforward generalization of Theorem $5.1$ of [2]. The main algorithmic difference is the usage of fractional matchings during the cut matching game and zeroing out in the random walk matrix rows/columns corresponding to vertices $v\notin\mathcal{T}$. (Recall that $\mathcal{T}=\{v\in V\mid\mu(v)>0\}$.)
@@
Given a graph $G=(V,E)$ of $m$ edges and a parameter $\phi$, there is a randomized algorithm that with high probability finds a partition of $V$ into clusters $V_{1},...,V_{k}$ such that $\forall i:\Phi^{\mu}_{G[V_{i}]}=\Omega(\phi)$ and $\sum_{i}{|E(V_{i},\bar{V_{i}})|}=O(\phi\mu(V)\log^{2}n)$. [^5] If we remove Assumption 2.1, the exact bound is $\sum_{i}{|E(V_{i},\bar{V_{i}})|}=O(\phi\mu(V)\log n\log\mu^{\star})$, where $\mu^{\star}=\frac{\max(\mu(\cdot))}{\min(\mu(\cdot))}$. The running time of the algorithm is $\tilde{O}(m)$.
The derivation of Theorem 3.2 using Theorem 3.1 uses the same methods as in [12, 7]. For completeness, we give this derivation in Appendix 5.
The rest of this section is organized as follows. In Subsection 3.1 we introduce a generalized version of the cut matching game of [2]. In Subsections 3.2 and 3.3 we present a cut player and a matching player, respectively. Finally, in Subsection 3.4 we present the algorithm for Theorem 3.2 using the cut and matching players.
@@ 3.1 Cut matching for $\mu$-expansion
The following is a generalized version of the cut matching games in [12, 2], suited for $\mu$.
The Cut-Matching game for $\mu$-expansion, with parameters $T$ and a vector $\mu$: The game is played on a series of graphs $G_{i}$. Initially, $G_{0}=\emptyset,A_{0}=V$. At iteration $t$, the cut player produces two weighted sets, $L_{t},R_{t}\subseteq A_{t}$, with weights $m_{v},\bar{m}_{v}$, respectively. The weights satisfy $m(L_{t})\leq\mu(A_{t})/8,\bar{m}(R_{t})\geq\mu(A_{t})/2$, and the total weight $m_{v}+\bar{m}_{v}$ of each $v\in V$ in $L_{t}\cup R_{t}$ is at most $\mu(v)$. The matching player responds with a set $S_{t}\subseteq A_{t}$ and a weighted matching $M_{t}$ that only matches vertices in $L_{t}\setminus S_{t}$ to vertices in $R_{t}\setminus S_{t}$. We set $G_{t+1}=G_{t}\cup M_{t},A_{t+1}=A_{t}\setminus S_{t}$. The union $G_{t}\cup M_{t}$ refers to the graph whose weighted adjacency matrix is the sum of the weighted adjacency matrices of $G_{t}$ and $M_{t}$. The game ends at iteration $T$, and the quality of the game is $r$, the near $\mu$-expansion of $A_{t}$ in $G_{t}$.
@@
    $I_{t}=diag(\mathbbm{1}_{A_{t}\cap\mathcal{T}})\in\mathop{\mathbb{R}}^{n\times n}$. Similarly, $\mathbbm{1}_{t}=I_{t}\cdot\mathbbm{1}$.
@@ 3.2 Cut player
Let $\delta=\Theta(\log{n})$ be a power of $2$ which we formally set later in the proof of Lemma 4.12.666We need $\delta$ to be a power of $2$ to apply Theorem  B.3 to a $\delta$ matrix power later on. The cut player implicitly maintains a $\mu$-stochastic flow matrix $F_{t}\in\mathop{\mathbb{R}}^{n\times n}$, and the graph $G_{t}$ which is the union of the matchings that it obtained so far from the matching player ($t$ is the index of the round). The flow $F_{t}$ and the graph $G_{t}$ have two crucial properties. First, we can embed $F_{t}$ in $G_{t}$ with $O(\frac{1}{\delta})$ congestion (Lemma 4.4) and embed $G_{t}$ in $G$ with $O(\frac{t}{\phi\log{n}})$ congestion (Lemma 4.5), see Definition 2.8. Second, after $T=\Theta(\log^{2}n)$ rounds, $A_{t}$ will be a near $(\Omega(1),\mu)$-expander in $F_{T}$ with high probability (Lemma 4.12). By Lemma 2.9, this implies that, with hight probability, $A_{t}$ will be a near $(\Omega(\phi),\mu)$-expander in $G$.
@@
We use the following lemma to partition (some of) the vertices of $A_{t}$ into two weighted sets $A_{t}^{l}$ and $A_{t}^{r}$. [^7] Note that this does not produce a bisection of $A_{t}$. The lemma is a continuous modification of Lemma 3.3 in [10]. For completeness, we include the proof in Appendix A.
Lemma 3.5 (Lemma 3.3 in [10]).
Given $u_{i}\in\mathop{\mathbb{R}}$ for all $i\in A_{t}$, such that $\sum_{i\in A_{t}}{\mu(i)u_{i}}=0$, we can find in time $O(|A_{t}|\cdot\log(|A_{t}|))$ a set of source nodes $A^{l}_{t}\subseteq A_{t}$, a set of target nodes $A^{r}_{t}\subseteq A_{t}$, weights $0\leq m_{i},\bar{m}_{i}\leq\mu(i)$ for $i\in A_{t}$, and a separation value $\eta$ such that:
@@
In turn, the matching player (see Subsection 3.3) returns a cut $(S_{t},A_{t}\setminus S_{t}$) and a fractional matching $M_{t}$ of total weight $m(A^{l}_{t}\setminus S_{t})$ between $A^{l}_{t}\setminus S_{t}$ and $A^{r}_{t}\setminus S_{t}$. Define $N_{t}=\frac{\delta-1}{\delta}U+\frac{1}{\delta}M_{t}$. The cut player then updates $F_{{t}}$ as follows: $F_{{t+1}}=N_{t}\cdot U^{-1}F_{t}U^{-1}N_{t}$. Intuitively, this applies lazy (with respect to $\delta$) averaging of the rows and columns of $F_{t}$ according to $M_{t}$. For the analysis, we update $G_{t+1}$ as $G_{t+1}=G_{t}\cup M_{t}$. [^8] We allow for parallel edges. Note that $G_{t+1}$ might include self-loops because of $A^{l}_{t}\cap A^{r}_{t}$.
@@ 3.3 Matching player
The matching player receives $A^{l}_{t}$ and $A^{r}_{t}$ and the current $A_{t}$. For a vertex $v\in V$, denote by $m_{v}$ the weight of $v$ in $A_{t}^{l}$, and by $\bar{m}_{v}$ the weight of $v$ in $A_{t}^{r}$ ($m_{v}+\bar{m}_{v}\leq\mu(v)$). The matching player solves the flow problem on $G[A_{t}]$, specified by the following lemma, that closely resembles Lemma $8.1$ of [7] when setting $V=A_{t}$. For completeness, we include the proof in Appendix A.2.
Lemma 3.6 (Lemma $8.1$, [7]).
Let $G=(V,E)$ be a graph with $n$ vertices and $m$ edges, let $A^{l},A^{r}\subseteq V$, and let $\phi>0$ be a parameter. For a vertex $v\in V$, denote by $m_{v}$ the weight of $v$ in $A^{l}$, and by $\bar{m}_{v}$ the weight of $v$ in $A^{r}$. Assume that $m_{v}+\bar{m}_{v}\leq\mu(v)$ and that $\bar{m}(A^{r})\geq\frac{1}{2}\mu(V),m(A^{l})\leq\frac{1}{8}\mu(V)$. We define the flow problem $\Pi(G)$, as the problem in which a source $s$ is connected to each vertex $v\in A^{l}$ with an edge of capacity $m_{v}$ and each vertex $v\in A^{r}$ is connected to a sink $t$ with an edge of capacity $\bar{m}_{v}$. Every edge of $G$ has the same capacity $c=\Theta\left(\frac{1}{\phi\log n}\right)$. A feasible flow for $\Pi(G)$ is a maximum flow that saturates all the edges outgoing from $s$. Then, in time $\tilde{O}(m)$, we can find either
@@
    A cut $S$ where $\Phi_{G}^{\mu}(S,V\setminus S)\leq\frac{7}{c}=O(\phi\log n)$, $\mu(V\setminus S)\geq\frac{1}{3}\mu(V)$ and a feasible flow for the problem $\Pi(G-S)$, where we only consider the sub-graph $G[V\setminus S\cup\{s,t\}]$ (that is, vertices $v\in A^{l}\setminus S$ are sources of $m_{v}$ units, and vertices $v\in A^{r}\setminus S$ are sinks of $\bar{m}_{v}$ units).
@@
Let $S_{t}$ be the cut returned by the lemma. If the lemma terminates with the first case, we set $S_{t}=\emptyset$. Using dynamic trees [11], we can decompose the returned flow into a set of (at most $m$) weighted paths, each carrying some flow from a vertex $u\in A_{t}^{l}\setminus S_{t}$ to a vertex $v\in A^{r}_{t}\setminus S_{t}$. If $u\in A^{l}_{t}\cap A^{r}_{t}$ then it is possible that a path starts and ends at $u$. Each $u\in A^{l}_{t}\setminus S_{t}$ is the starting point of exactly $m_{u}\leq\mu(u)$ flow, and each $v\in A^{r}_{t}\setminus S_{t}$ is the endpoint of at most $\bar{m}_{v}\leq\mu(v)$ flow. Define the weighted matching $\tilde{M_{t}}$ as $\tilde{M_{t}}=((u_{i},v_{i},w_{i}))_{i=1}^{|A^{l}_{t}\setminus S_{t}|}$, where $u_{i}$ and $v_{i}$ are the endpoints of path $i$ which carries $w_{i}$ flow. We can view $\tilde{M_{t}}$ as a symmetric $n\times n$ matrix, such that $\tilde{M_{t}}(u,v)=\tilde{M_{t}}(v,u)=\tilde{w}_{uv}$, where $\tilde{w}_{uv}$ is the total flow sent on paths between $u$ and $v$. We turn $\tilde{M_{t}}$ into a $\mu$-stochastic matrix by increasing its diagonal entries by $\mu-\tilde{M_{t}}\mathbbm{1}_{n}$. Formally, we set $M_{t}\mathrel{\mathop{:}}=\tilde{M_{t}}+\operatorname{\mathbf{diag}}(\mu-\tilde{M_{t}}\mathbbm{1}_{n})$. Notice that $\mu-\tilde{M_{t}}\mathbbm{1}_{n}$ has only non-negative entries, so $M_{t}$ also has non-negative entries. The matching player sends back $M_{t},S_{t}$ as the response to the subsets $A^{l}_{t}$ and $A^{r}_{t}$ given by the cut player.

### v2

@@ 3 Balanced Sparse-Cut via spectral Cut-Matching
In this section we state our main theorem and present our algorithm, whose analysis appears in the next section. The result of [agassy2022expander] for finding a balance sparse cut for the case of conductance is a special case of our theorem in which the vertex measure corresponds to degrees (i.e. $\mu(v)=\deg_{G}(v)$ for every $v\in V$).
@@
Given a graph $G=(V,E)$ with $n$ vertices and $m$ edges, a vertex measure $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$ that satisfies Assumption 2.1 and a parameter $\phi>0$, there exists a randomized algorithm which takes $\tilde{O}\left(m\right)$ time [^5] The $\tilde{O}$ hides $\log n$ factors . and must end with one of the following three cases:
@@
    We certify that $G$ has $\mu$-expansion $\Phi^{\mu}(G)=\Omega(\phi)$.
@@
    We find a cut $(R,A)$ in $G$ of $\mu$-expansion $\Phi^{\mu}_{G}(R,A)=O(\phi\log n)$, and $\mu(R),\mu(A)$ are both $\Omega\!\left(\frac{\mu(V)}{\log n}\right)$, i.e, we find a relatively balanced low conductance cut.
@@
    We find a cut $(R,A)$ with $\Phi^{\mu}_{G}(R,A)\leq c_{0}\phi\log n$ for some constant $c_{0}>0$, $0<\mu(R)\leq\frac{\mu(V)}{10c_{0}\log n}$, and $A$ is a near $(\Omega(\phi),\mu)$-expander in $G$.
The correctness of the algorithm holds with high probability. [^6] By “with high probability” we mean a probability of at least $1-\frac{1}{n^{\ell}}$, where $\ell$ is an arbitrary constant that affects our running time linearly..
This theorem is a straightforward generalization of [agassy2022expander, Theorem 5.1]. The main algorithmic difference is the usage of fractional matchings during the cut matching game and zeroing out in the random walk matrix rows/columns corresponding to vertices $v\notin\mathcal{T}$. (Recall that $\mathcal{T}=\{v\in V\mid\mu(v)>0\}$.)
@@
Given a graph $G=(V,E)$ with $n$ vertices and $m$ edges, a vertex measure $\mu\in\mathop{\mathbb{R}}^{n}_{\geq 0}$ that satisfies Assumption 2.1, and a parameter $\phi>0$, there is a randomized algorithm that with high probability finds a partition of $V$ into clusters $V_{1},...,V_{k}$ such that $\forall i:\Phi^{\mu}_{G[V_{i}]}=\Omega(\phi)$ and $\sum_{i}{|E(V_{i},\overline{V_{i}})|}=O(\phi\mu(V)\log^{2}n)$. The running time of the algorithm is $\tilde{O}(m)$.
The derivation of Theorem 3.2 using Theorem 3.1 uses the same methods as in [saranurak2019expander, LNPSsoda13, agassy2022expander]. For completeness, we give this derivation in Section 5.
The rest of this section is organized as follows. In Section 3.1 we introduce a generalized version of the cut matching game of [agassy2022expander]. In Sections 3.2 and 3.3 we present a cut player and a matching player, respectively. Finally, in Section 3.4 we present the algorithm for Theorem 3.2 using the cut and matching players.
@@ 3.1 Cut matching for $\mu$-expansion
The following is a generalized version of the cut matching games in [saranurak2019expander, agassy2022expander], suited for $\mu$.
The Cut-Matching game for $\mu$-expansion, with parameters $T$ and a vector $\mu$: The game is played on a series of graphs $G_{i}$. Initially, $G_{0}=\emptyset,A_{0}=V$. At iteration $t$, the cut player produces two sets, $L_{t},R_{t}\subseteq A_{t}$, and weights $m_{t}:L_{t}\to\mathop{\mathbb{R}}_{\geq 0},\bar{m}_{t}:R_{t}\to\mathop{\mathbb{R}}_{\geq 0}$. [^7] We denote $m_{t}(v)=0$ for each $v\notin L_{t}$ and $\bar{m}_{t}(v)=0$ for each $v\notin R_{t}$. The weights satisfy $m_{t}(L_{t})\leq\frac{\mu(A_{t})}{8},\bar{m}_{t}(R_{t})\geq\frac{\mu(A_{t})}{2}$, and the total weight $m_{t}(v)+\bar{m}_{t}(v)$ of each $v\in L_{t}\cup R_{t}$ is at most $\mu(v)$. The matching player responds with a set $S_{t}\subseteq A_{t}$ and a $\mu$-matching $M_{t}$ that only matches vertices in $L_{t}\setminus S_{t}$ to vertices in $R_{t}\setminus S_{t}$. We set $G_{t+1}=G_{t}\cup M_{t},A_{t+1}=A_{t}\setminus S_{t}$. The union $G_{t}\cup M_{t}$ refers to the graph whose weighted adjacency matrix is the sum of the weighted adjacency matrices of $G_{t}$ and $M_{t}$. The game ends at iteration $T$, and the quality of the game is $r$, the near $\mu$-expansion of $A_{t}$ in $G_{t}$.
@@
    $I_{t}=\operatorname{\mathbf{diag}}(\mathbbm{1}_{A_{t}\cap\mathcal{T}})\in\mathop{\mathbb{R}}^{n\times n}$. Similarly, $\mathbbm{1}_{t}=I_{t}\cdot\mathbbm{1}$.
@@ 3.2 Cut player
Let $\delta=\Theta(\log{n})$ be a power of $2$ which we formally set later in the proof of Lemma 4.12.888We need $\delta$ to be a power of $2$ to apply Theorem B.3 to a $\delta$ matrix power later on. The cut player implicitly maintains a $\mu$-stochastic matrix $F_{t}\in\mathop{\mathbb{R}}^{n\times n}$, and the graph $G_{t}$ which is the union of the matchings that it obtained so far from the matching player ($t$ is the index of the round). The matrix $F_{t}$ and the graph $G_{t}$ have two crucial properties. First, we can embed (recall Definition 2.7) $F_{t}$ in $G_{t}$ with $O(\frac{1}{\delta})$ congestion (Lemma 4.4) and embed $G_{t}$ in $G$ with $O(\frac{t}{\phi\log{n}})$ congestion (Lemma 4.5). Second, after $T=\Theta(\log^{2}n)$ rounds, $A_{t}$ will be a near $(\Omega(1),\mu)$-expander in $F_{T}$ with high probability (Lemma 4.12). By Lemma 2.8, this implies that, with high probability, $A_{t}$ will be a near $(\Omega(\phi),\mu)$-expander in $G$.
@@
We use the following lemma to partition (some of) the vertices of $A_{t}$ into two weighted sets $A_{t}^{l}$ and $A_{t}^{r}$. [^9] Note that this does not produce a bisection of $A_{t}$. The lemma is a continuous modification of [racke2014computing, Lemma 3.3]. For completeness, we include the proof in Appendix A.
Lemma 3.5 ([racke2014computing, Lemma 3.3]).
Given $u_{i}\in\mathop{\mathbb{R}}$ for all $i\in A_{t}$, such that $\sum_{i\in A_{t}}{\mu(i)u_{i}}=0$, we can find in time $O(|A_{t}|\cdot\log(|A_{t}|))$ a set of source nodes $A^{l}_{t}\subseteq A_{t}$, a set of target nodes $A^{r}_{t}\subseteq A_{t}$, weights $0\leq m_{t}(i)\leq\mu(i)$ for all $i\in A^{l}_{t}$, $0\leq\bar{m}_{t}(i)\leq\mu(i)$ for all $i\in A^{r}_{t}$, and a separation value $\eta\in\mathop{\mathbb{R}}$ such that:
@@
In turn, the matching player (see Subsection 3.3) returns a cut $(S_{t},A_{t}\setminus S_{t}$) and a fractional matching $M_{t}$ of total weight $m_{t}(A^{l}_{t}\setminus S_{t})$ between $A^{l}_{t}\setminus S_{t}$ and $A^{r}_{t}\setminus S_{t}$. Define $N_{t}=\frac{\delta-1}{\delta}U+\frac{1}{\delta}M_{t}$, the lazy random walk that takes a step according to $M_{t}$ with probability $\frac{1}{\delta}$, and otherwise takes a step according to $U$. The cut player then updates $F_{{t}}$ as follows: $F_{{t+1}}=N_{t}\cdot U^{-1}F_{t}U^{-1}N_{t}$. Intuitively, this applies a weighted average of the rows and columns of $F_{t}$ according to $M_{t}$, where each row/column “keeps” $\frac{\delta-1}{\delta}$ fraction of its value. For the analysis, we update $G_{t+1}$ as $G_{t+1}=G_{t}\cup M_{t}$. [^10] We allow for parallel edges. Note that $G_{t+1}$ might include self-loops because of vertices in $A^{l}_{t}\cap A^{r}_{t}$.
@@ 3.3 Matching player
The matching player receives $A^{l}_{t}$ and $A^{r}_{t}$ and the current $A_{t}$. For a vertex $v\in V$, denote by $m_{t}(v)$ the weight of $v$ in $A_{t}^{l}$, and by $\bar{m}_{t}(v)$ the weight of $v$ in $A_{t}^{r}$ ($m_{t}(v)+\bar{m}_{t}(v)\leq\mu(v)$). The matching player solves the flow problem on $G[A_{t}]$, specified by the following lemma, that closely resembles [LNPSsoda13, Lemma 8.1] when setting $V=A_{t}$. For completeness, we include the proof in Appendix A.2.
Lemma 3.6 ([LNPSsoda13, Lemma 8.1]).
Let $G=(V,E)$ be a graph with $n$ vertices, let $A^{l},A^{r}\subseteq V$, and let $\phi>0$ be a parameter. For a vertex $v\in V$, denote by $m(v)$ the weight of $v$ in $A^{l}$, and by $\bar{m}(v)$ the weight of $v$ in $A^{r}$. Assume that $m(v)+\bar{m}(v)\leq\mu(v)$ and that $\bar{m}(A^{r})\geq\frac{1}{2}\mu(V),m(A^{l})\leq\frac{1}{8}\mu(V)$. We define the flow problem $\Pi(G)$, as the problem in which a source $s$ is connected to each vertex $v\in A^{l}$ with an edge of capacity $m(v)$ and each vertex $v\in A^{r}$ is connected to a sink $t$ with an edge of capacity $\bar{m}(v)$. Every edge of $G$ has the same capacity $c=\Theta\left(\frac{1}{\phi\log n}\right)$. A feasible flow for $\Pi(G)$ is a maximum flow that saturates all the edges outgoing from $s$. Then, there exists a randomized algorithm, which in time $\tilde{O}(|E|)$, finds either
@@
    A cut $S$ where $\Phi_{G}^{\mu}(S,V\setminus S)\leq\frac{7}{c}=O(\phi\log n)$, $\mu(V\setminus S)\geq\frac{1}{3}\mu(V)$ and a feasible flow for the problem $\Pi(G-S)$, where we only consider the sub-graph $G[V\setminus S\cup\{s,t\}]$ (that is, vertices $v\in A^{l}\setminus S$ are sources of $m(v)$ units, and vertices $v\in A^{r}\setminus S$ are sinks of $\bar{m}(v)$ units).
The correctness of the algorithm holds with high probability.
@@
Let $S_{t}$ be the cut returned by the lemma. If the lemma terminates with the first case, we set $S_{t}=\emptyset$. Using dynamic trees [ST83], we can decompose the returned flow into a set of (at most $m$) weighted paths, each carrying some flow from a vertex $u\in A_{t}^{l}\setminus S_{t}$ to a vertex $v\in A^{r}_{t}\setminus S_{t}$. If $u\in A^{l}_{t}\cap A^{r}_{t}$ then it is possible that a path starts and ends at $u$. Each $u\in A^{l}_{t}\setminus S_{t}$ is the starting point of exactly $m_{t}(u)\leq\mu(u)$ units of flow, and each $v\in A^{r}_{t}\setminus S_{t}$ is the target of at most $\bar{m}_{t}(v)\leq\mu(v)$ units of flow. Define the weighted matching $\tilde{M_{t}}$ as $\tilde{M_{t}}=((u_{i},v_{i},w_{i}))_{i=1}^{|A^{l}_{t}\setminus S_{t}|}$, where $u_{i}$ and $v_{i}$ are the endpoints of path $i$ which carries $w_{i}$ flow. We can view $\tilde{M_{t}}$ as a symmetric $n\times n$ matrix, such that $\tilde{M_{t}}(u,v)=\tilde{M_{t}}(v,u)=w_{uv}$, where $w_{uv}$ is the total flow sent on paths between $u$ and $v$. We turn $\tilde{M_{t}}$ into a $\mu$-stochastic matrix by increasing its diagonal entries by $\mu-\tilde{M_{t}}\mathbbm{1}_{n}$. Formally, we set $M_{t}\mathrel{\mathop{:}}=\tilde{M_{t}}+\operatorname{\mathbf{diag}}(\mu-\tilde{M_{t}}\mathbbm{1}_{n})$. Notice that $\mu-\tilde{M_{t}}\mathbbm{1}_{n}$ has only non-negative entries, so $M_{t}$ also has non-negative entries. The matching player sends back $M_{t},S_{t}$ as the response to the subsets $A^{l}_{t}$ and $A^{r}_{t}$ given by the cut player.